[Brahms-l] Shift report 20040131 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Sat Jan 31 2004 - 00:01:46 EST
------------------------------------------------------------
Date: Sat Jan 31 00:01:45 2004, Shift: 16:00-24:00
Supervisor: Ionut ARSENE, with: 
Log book#15 pages 46 - 46

Shift Summary:
------------------------------------------------------------
1. Beam for 3.5 hours

2. Settings:
 MRS@40deg with D5@1050
 FS@3deg with D1@1723 but the trigger6 and trig2 was off.
  Runs:
 9970 - 200k trig3
 9971 - 180k trig3
 9972 - 370k trig3

3. No trigger 2 and 6 for these runs. FS was moved to 3
degrees. We asked for an access (granted), and solved the
FS problem (thanks to Hiro). 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jan 31 00:02:03 2004

This archive was generated by hypermail 2.1.8 : Sat Jan 31 2004 - 00:02:24 EST